OpenStudy (anonymous):

This time rectangle R has varying length l and width w but with a constant area of 4 square feet. a) Express the perimeter P as a function of length l. What type of function is P? What is the domain of P? b) Describe the asymptotic behavior of P. What can you say about rectangle Rbecause of this behavior? Could you have made a similar statement about Rback in Task 1? c) For what values of land w will the perimeter of R be the least? Give a geometric explanation. Be sure to include a graph with relevant points labeled.

OpenStudy (anonymous):

a) Write formula for the perimeter in terms of l and w

OpenStudy (anonymous):

How do you do that? I'm really confused

OpenStudy (anonymous):

 Perimeter = total length of the boundary. Now write the formula for perimeter.

OpenStudy (anonymous):

Like p= 4 times ( L + W)?

OpenStudy (anonymous):

p = l+l+w+w = 2l+2w = 2(l+w) So, p = 2times (l+w)

OpenStudy (anonymous):

Now area = const Area of a rectangle = l*w so l*w = 4 ft^2 w = 4/l units

OpenStudy (anonymous):

Put this w in the formula for perimeter you'll get p as function of l

OpenStudy (anonymous):

Is all of this for section a ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So how do we find the domain?

OpenStudy (anonymous):

What is the form of p in l ?

OpenStudy (anonymous):

I got p= 2?

OpenStudy (anonymous):

Since p = 2*(l+w) and w = 4/l ,so p = 2*(l +4/l)

OpenStudy (anonymous):

Check what values l can hold for a meaningful p.

OpenStudy (anonymous):

This is where I'm at .. How do I find the domain?

OpenStudy (anonymous):

Think. What if l =0 ?

OpenStudy (anonymous):

Well if L was zero everything would equal zero wouldn't it?

OpenStudy (anonymous):

No. check it again.

OpenStudy (anonymous):

Am I plugging it to p= 2 (l+4/l)?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

But 4/l is undefined because you can't divide 4/0

OpenStudy (anonymous):

Right! Good going. Length of a rectangle being zero has no meaning. Can you think about any other value of l which will have no meaning either logically or plugging it in the formula for p

OpenStudy (anonymous):

Or would my Answer be 8??

OpenStudy (anonymous):

Don't mind my last message haha I was just overthinking things

OpenStudy (anonymous):

4/0 is zero is undefined. so p is undefined. Could think of any other l for which p is undefined

OpenStudy (anonymous):

So is my domain that? That's it's undefined ?

OpenStudy (anonymous):

Nope

OpenStudy (anonymous):

Domain means values of l for which p is defined.

OpenStudy (anonymous):

So whatnot my domain?

OpenStudy (anonymous):

What is my domain? Zero?

OpenStudy (anonymous):

Ok. l can take any value except 0. So the domain is all real numbers except zero. You can write it like this \[l \in R - \left\{ 0 \right\}\]

OpenStudy (anonymous):

ohh okay thank you! and What can you say about rectangle R because of this behavior?

OpenStudy (anonymous):

For what values of l and w will the perimeter of R be the least? Give a geometric explanation. Be sure to include a graph with relevant points labeled how do I graph with relevant points labeled?

OpenStudy (anonymous):

what type of function is P?